OpenStudy (nikkibxox):

Checking for Extranuous Solutions Help Among Other Things. Will give Medals! 1.) Solve 9|9-8x|=2x+3 and check for extranous solutions 2.) Which equation defines the graph of y=x^2 after it is shifted vertically 3 units up and horizontally 2 units right? a) y=(x+2)^2+3 b) y=(x-3)^2+2 c) y=(x-2)^2+3 d) y=(x+3)^2+2 3.) A candle is 17 in tall after burning for 3 hours. After 5 hours it is 15 inches tall. Write a linear equation to model the relationship between height (h) and the time(t). Predict the height of the candle at 8 hours. Show all work and Explain, Please! Thank you!

OpenStudy (nikkibxox):

@texaschic101

OpenStudy (nikkibxox):

@carlyleukhardt

OpenStudy (logan13):

ill help u if you help me

OpenStudy (nikkibxox):

@dan815

OpenStudy (nikkibxox):

@Chillsquad

OpenStudy (owlcoffee):

For number (1), we have to take in consideration the very definition of Absolute value: \[9\left| 9-8x \right|=2x+3\] And the definition of absolute value is: \(\left| \alpha \right|=\) \(\alpha , x \ge 0\) \(-\alpha , x<0\) So, therefore, we have to consider two scenarios for the solution of the equation: \[(1) \rightarrow 9(9-8x)=2x+3\] \[(2) \rightarrow (-9)(9-8x)=2x+3\] And you solve them both independantly to get the solutions.

OpenStudy (nikkibxox):

Thanks!!

OpenStudy (danjs):

do you know what the extraneous solutions could be if any